Session: Feb 2002 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (X.Y. Huang at Mon Feb 18 06:10:09 MET 2002) Success

Medicina (Carlo Migoni at Mon Feb 18 11:13:59 MET 2002) Success

Jodrell_Bank (Alastair Gunn at Mon Feb 18 11:30:22 MET 2002) Minor failures
Bad count readback in VC04, therefore Tsys corrupted. Voltage still looks okay on VC04 so data should be good. No other known problems.

Cambridge (Alastair Gunn at Mon Feb 18 11:30:42 MET 2002) Failure
Cambridge did not observe due to serious fault with recorder.

Effelsberg (A. Kraus at Mon Feb 18 12:01:24 MET 2002) Success

Torun (Kazimierz Borkowski at Mon Feb 18 13:36:52 MET 2002) Success
No known problems (Tsys cal working, the offending BBC7 not used).

Onsala (John Conway at Tue Feb 19 08:40:23 MET 2002) Failure
Did not observe dues to failure of Capstan Motor EVN spare is being shipped and should be installed by Tuesday afternoon

Westerbork (Tony Foley at Tue Feb 19 14:40:39 MET 2002) Minor failures
scans at 22:50 and at 00:05 out of HA range. One telescope out of array (RTC was jumping in phase)

Urumqi (Liu Xiang at Tue Feb 19 14:40:39 MET 2002) Minor failures
Before Feb.17, 17:00, #13 H-maser was unlocked, so changed using #11
H-maser, but the later one was frequency unstable-having shift, caused
the H-maser to GPS and the Formatter 's error was increased (by less
than 0.5 us). On UT0:00 little problem with date, resolved in a short time.